OpenStudy (anonymous):

PLEASE HELP?????Consider the following equation -4/x-4 +2/x+8=1 state any restrictions on the variable.

OpenStudy (anonymous):

\[\frac{-4}{x-4}+\frac{2}{x+8}=1\] no denominator can be \(0\) \(x\) can neither be \(4\) nor \(-8\)

OpenStudy (anonymous):

will it be nor restrictions